### v3

@@ 3.3 Solutions of Issues (RQ3)
The taxonomy of solutions for microservices issues is provided in Table 4. It is worth mentioning that not all the issue discussions provide the information about their solutions. Therefore, we identified 1,899 issue discussions containing the information about the solutions. The taxonomy of solutions is derived by mining developer discussions (i.e., 1,899 solutions), conducting practitioner interviews (i.e., 36 solutions, see Section 2.2.2), and conducting a survey (i.e., one instance of solution, see Section 3.4). Hence, we got a total of 1,936 solutions. We identified a total of 196 types of solutions that can be classified in 8 categories and 35 subcategories. Due to space limitations, we only list the top two types of solutions for each subcategory in Table 4. The detail of the types of solutions can be found in the dataset [19]. The results show that Fix Artifacts (1056 out of 1936), Add Artifacts (360 out of 1936), and Modify Artifacts (210 out of 1936) are the top three categories of solutions. Each solution category is briefly discussed below.
1. Fix Artifacts (1056/1936, 54.54%): During the analysis of developer discussions about microservices issues, we identified that most developers did not explicitly mention any solution to the problems. They fixed the issue in the local repository and send the fixed code (e.g., through a pull request) to the maintainer of the public repository. In this case, from the developer discussions, we could not find exactly what they added, removed, or modified in the project to fix a specific issue. Therefore, we named this category Fix Artifacts. We identified and classified 25 types of fix artifacts solutions in 4 subcategories (see Table 4 and the Solution Taxonomy sheet in [19]). Each of them is briefly described below.
@@
We identified and classified 33 types of solutions in 10 subcategories (see Table 4 and the Solution Taxonomy sheet in [19]). Each of them is briefly described below.
  • Add Features and Services (131, 6.76%): Developers added required features and services to microservices systems. System features and services refer to a process that accepts one or more inputs and returns outputs for particular system functionality [35]. We identified 5 types of solutions in this subcategory, and among them, add missing features, add security features, and add communication protocols are the top three types of solutions.
@@
3. Modify Artifacts (212/1936, 10.95%): Besides adding new artifacts to the existing system to address the microservices issues, we also identified a large number of solutions in which the developers explicitly mentioned how they modified modules, services, packages, APIs, scripts, methods, objects, data types, identifiers, databases, and documentation to address the microservices issues. We identified and classified 31 types of solutions in 5 subcategories (see Table 4 and the Solution Taxonomy sheet in [19]). Each of them is briefly described below.
@@
4. Remove Artifacts (39/1936, 2.01%): This category covers the solutions in which artifacts are removed to address several types of microservices issues. We identified and classified 14 types of solutions in 5 subcategories (see Table 4 and the Solution Taxonomy sheet in [19]). Each of them is briefly described below.
@@
5. Manage Infrastructure (160/1936, 8.26%): This category captures the solutions that are based on efficient resource utilization for addressing the microservices issues. We identified and classified 20 types of solutions in 2 subcategories (see Table 4 and the Solution Taxonomy sheet in [19]). Each of them is briefly described below.
@@
6. Manage Configuration and Execution (50/1936, 2.58%): Managing configuration and execution enables developers to track changes in microservices systems and their consuming applications over time, for example, the ability to track the version history of configuration changes for multiple instances of microservices systems. We identified and classified 13 types of solutions in 2 subcategories (see Table 4 and the Solution Taxonomy sheet in [19]). Each of them is briefly described below.
@@
We identified and classified 21 types of solutions in two subcategories (see Table 4 and the Solution Taxonomy sheet in [19]). Each of them is briefly described below.
@@
8. Import/Export Artifacts (12/1936, 0.61%): We found several issues that can be fixed by importing and exporting various artifacts. We identified and classified 3 types of solutions in 2 subcategories (see Table 4 and the Solution Taxonomy sheet in [19]). Each of them is briefly described below.
@@ 3.4 Practitioners’ Perspective (RQ4)
[…]
Practitioners’ Perspective on Issues: We asked the microservices participants which issue they faced while developing microservices systems. The majority of the respondents mentioned that they face all of the issues while developing microservices systems. The practitioners’ responses presented in Table 5 show that IC1: Technical Debt (46.67% Very Often, 24.00% Often), IC2: Continuous Integration and Delivery Issue (26.67% Very Often, 42.67% Often), and IC5: Security Issue (18.67% Very Often, 64.00% Often) occur most frequently than other categories of issues. Some practitioners also suggested 6 types of issues (with 9 instances of issues) that were not part of the initial taxonomy of issues in microservices systems. We added microservices practitioners’ suggested types of issues in Figure 4 and the Issue Taxonomy sheet in [19]. One representative quotation about other types of issues is depicted below.
[…]
@@
Practitioners’ Perspective on Causes: Our survey participants also confirmed the identified causes that could lead to issues in microservices systems. The practitioners’ responses presented in Table 5 show that CC1: General Programming Error (41.33% Strongly Agree, 38.67% Agree), CC8: Fragile Code (23.33% Strongly Agree, 46.67% Agree), and CC2: Missing Features and Artifacts (12.67% Strongly Agree, 55.33% Agree) are the top three cause categories. Some practitioners also suggested 5 types of causes (with 11 instances of causes) that were not part of the initial taxonomy of causes in microservices systems. We added microservices practitioners’ suggested types of causes in Table 5 and the Cause Taxonomy sheet in [19]. One representative quotation about suggested types of causes is depicted below.
[…]
@@
Practitioners’ Perspective on Solutions: We also recorded the practitioners’ perspective about the solutions for the issues occurring during microservices system development. The practitioners’ responses presented in Table 5 show that SC1: Add Artifacts (42.00% Strongly Agree, 39.33% Agree), SC3: Modify Artifact (27.33% Strongly Agree, 37.33% Agree), and SC6: Manage Configuration and Execution (18.00% Strongly Agree, 48.67% Agree) and are the top three solution categories that have been used to address the microservices issues. One practitioner also suggested one type of solution (with one instance of solution) that was not part of the initial taxonomy of solutions in microservices systems. We added microservices practitioners’ suggested types of solutions in Table 5 and the Solution Taxonomy sheet in [19]. One representative quotation about the suggested solutions is depicted below.
[…]
@@
[…]
@@
4 Discussion
We analyze in this section the relationship between the issues, causes, and solutions (Section 4.1) followed by presenting the implications of the research (Section 4.2).
@@ 4.1 Relationship between issues, causes, solutions
While the taxonomy in Figure 4 provides a categorization of issues in microservices systems, a mapping between the issues, their causes, and solutions is presented in Figure 5. Mapping diagrams are frequently used in systematic mapping studies—relying on bubble plots—to correlate data or concepts along different dimensions [36]. We have chosen the mapping diagram to depict the issues on the Y-axis and their causes on the X-axis, and to present the solutions that can address these issues at the intersections of the X and Y axes. Eight different types of symbols, labeled at the top of the figure, are used to represent various types of solutions. The interpretation of the mapping in Figure 5 involves locating a given issue on the Y-axis, associating this issue with its cause(s) on the X-axis, and identifying the possible solutions to address the issue, as elaborated and exemplified below. Figure 5 indicate that it is possible that one issue could cause other issues. Similarly, one cause may be the reason for several issues, or several causes may contribute to one issue.
@@
The mapping in Figure 5 can have a diverse interpretation based on the intent of the analysis that may include but is not limited to frequency analysis and data correlations. It is virtually impossible to elaborate on all possible interpretations, however; to exemplify some of the possible interpretations can be as follows:
  • What are the most and least frequently occurring issues in microservices systems? As per the mapping, the most frequent (top 3) issues are related to Technical Debt, CI/CD, and Exception Handling, representing a total of 764 identified issues. On the other hand, the least frequent issues relate to Organizational, Update and Installation, and Typecasting issues.
  • What are the most and least common causes of a specific category of issues? The mapping of the causes suggests that General Programming Errors, Invalid Configuration and Communication Problems along with Legacy Versions, Compatibility and Dependency Problem are the most common causes of the issues. Similarly, Fragile Code represents the least common cause for issues in microservices systems.
  • What are the most and least recurring solutions to fix a specific category of issues? Fixing Artifacts, Add Artifacts, and Modify Artifacts represent the most recurring solutions to address microservices issues, whereas importing/Export Artifacts, Upgrade Tools and Platforms, and Manage Configuration and Execution categories represent the least recurring solutions to address microservices issues.

### v4

@@ 3.3 Solutions of Issues (RQ3)
The taxonomy of solutions for microservices issues is provided in Table 4. It is worth mentioning that not all the issue discussions provide the information about their solutions. Therefore, we identified 1,899 issue discussions containing the information about the solutions. The taxonomy of solutions is derived by mining developer discussions (i.e., 1,899 solutions), conducting practitioner interviews (i.e., 36 solutions, see Section 2.2.2), and conducting a survey (i.e., one instance of solution, see Section 3.4). Hence, we got a total of 1,936 solutions. We identified a total of 196 types of solutions that can be classified in 8 categories and 35 subcategories. Due to space limitations, we only list the top two types of solutions for each subcategory in Table 4. The detail of the types of solutions can be found in the dataset [replpack]. The results show that Fix Artifacts (1056 out of 1936), Add Artifacts (360 out of 1936), and Modify Artifacts (210 out of 1936) are the top three categories of solutions. Each solution category is briefly discussed below.
1. Fix Artifacts (1056/1936, 54.54%): During the analysis of developer discussions about microservices issues, we identified that most developers did not explicitly mention any solution to the problems. They fixed the issue in the local repository and send the fixed code (e.g., through a pull request) to the maintainer of the public repository. In this case, from the developer discussions, we could not find exactly what they added, removed, or modified in the project to fix a specific issue. Therefore, we named this category Fix Artifacts. We identified and classified 25 types of fix artifacts solutions in 4 subcategories (see Table 4 and the Solution Taxonomy sheet in [replpack]). Each of them is briefly described below.
@@
We identified and classified 33 types of solutions in 10 subcategories (see Table 4 and the Solution Taxonomy sheet in [replpack]). Each of them is briefly described below.
  • Add Features and Services (131, 6.76%): Developers added required features and services to microservices systems. System features and services refer to a process that accepts one or more inputs and returns outputs for particular system functionality [paolucci2002semantic]. We identified 5 types of solutions in this subcategory, and among them, add missing features, add security features, and add communication protocols are the top three types of solutions.
@@
3. Modify Artifacts (212/1936, 10.95%): Besides adding new artifacts to the existing system to address the microservices issues, we also identified a large number of solutions in which the developers explicitly mentioned how they modified modules, services, packages, APIs, scripts, methods, objects, data types, identifiers, databases, and documentation to address the microservices issues. We identified and classified 31 types of solutions in 5 subcategories (see Table 4 and the Solution Taxonomy sheet in [replpack]). Each of them is briefly described below.
@@
4. Remove Artifacts (39/1936, 2.01%): This category covers the solutions in which artifacts are removed to address several types of microservices issues. We identified and classified 14 types of solutions in 5 subcategories (see Table 4 and the Solution Taxonomy sheet in [replpack]). Each of them is briefly described below.
@@
5. Manage Infrastructure (160/1936, 8.26%): This category captures the solutions that are based on efficient resource utilization for addressing the microservices issues. We identified and classified 20 types of solutions in 2 subcategories (see Table 4 and the Solution Taxonomy sheet in [replpack]). Each of them is briefly described below.
@@
6. Manage Configuration and Execution (50/1936, 2.58%): Managing configuration and execution enables developers to track changes in microservices systems and their consuming applications over time, for example, the ability to track the version history of configuration changes for multiple instances of microservices systems. We identified and classified 13 types of solutions in 2 subcategories (see Table 4 and the Solution Taxonomy sheet in [replpack]). Each of them is briefly described below.
@@
We identified and classified 21 types of solutions in two subcategories (see Table 4 and the Solution Taxonomy sheet in [replpack]). Each of them is briefly described below.
@@
8. Import/Export Artifacts (12/1936, 0.61%): We found several issues that can be fixed by importing and exporting various artifacts. We identified and classified 3 types of solutions in 2 subcategories (see Table 4 and the Solution Taxonomy sheet in [replpack]). Each of them is briefly described below.
@@ 3.4 Practitioners’ Perspective (RQ4)
We conducted a cross-sectional survey to evaluate the taxonomies of issues, causes, and solutions in microservices systems built in RQ1, RQ2, and RQ3. We provided a list of 19 issue categories and asked survey participants to respond to each category on a 5-point Likert scale (Very Often, Often, Sometimes, Rarely, Never). Similarly, regarding causes and solutions, we provided 8 cause categories and 8 solution categories, and asked practitioners to respond to each category on a 5-point Likert scale (Strongly Agree, Agree, Neutral, Disagree, Strongly Disagree). We also asked three open-ended questions to identify the missing issues, causes, and solutions in the provided categories. We received 150 valid responses completed by microservices practitioners from 42 countries across 6 continents. The results of RQ4 are summarized in two tables (i.e., Table 5 and Table 6). This section also provides representative quotations from practitioners for answering open-ended questions with the sign. The practitioners’ perspectives on the issues, causes, and solutions categories in microservices systems are presented in Table 5. Due to the limited space, we only presented the ‘percentage’ and ‘mean’ values of the practitioners’ responses to each category of issues, causes, and solutions. The survey results about the practitioners’ perspectives on microservices issues, causes, and solutions are briefly reported below.
Practitioners’ Perspective on Issues: We asked the microservices participants which issue they faced while developing microservices systems. The majority of the respondents mentioned that they face all of the issues while developing microservices systems. The practitioners’ responses presented in Table 5 show that IC1: Technical Debt (46.67% Very Often, 24.00% Often), IC2: Continuous Integration and Delivery Issue (26.67% Very Often, 42.67% Often), and IC5: Security Issue (18.67% Very Often, 64.00% Often) occur most frequently than other categories of issues. Some practitioners also suggested 6 types of issues (with 9 instances of issues) that were not part of the initial taxonomy of issues in microservices systems. We added microservices practitioners’ suggested types of issues in Figure 4 and the Issue Taxonomy sheet in [replpack]. One representative quotation about other types of issues is depicted below.
@@
Practitioners’ Perspective on Causes: Our survey participants also confirmed the identified causes that could lead to issues in microservices systems. The practitioners’ responses presented in Table 5 show that CC1: General Programming Error (41.33% Strongly Agree, 38.67% Agree), CC8: Fragile Code (23.33% Strongly Agree, 46.67% Agree), and CC2: Missing Features and Artifacts (12.67% Strongly Agree, 55.33% Agree) are the top three cause categories. Some practitioners also suggested 5 types of causes (with 11 instances of causes) that were not part of the initial taxonomy of causes in microservices systems. We added microservices practitioners’ suggested types of causes in Table 5 and the Cause Taxonomy sheet in [replpack]. One representative quotation about suggested types of causes is depicted below.
@@
Practitioners’ Perspective on Solutions: We also recorded the practitioners’ perspective about the solutions for the issues occurring during microservices system development. The practitioners’ responses presented in Table 5 show that SC1: Add Artifacts (42.00% Strongly Agree, 39.33% Agree), SC3: Modify Artifact (27.33% Strongly Agree, 37.33% Agree), and SC6: Manage Configuration and Execution (18.00% Strongly Agree, 48.67% Agree) and are the top three solution categories that have been used to address the microservices issues. One practitioner also suggested one type of solution (with one instance of solution) that was not part of the initial taxonomy of solutions in microservices systems. We added microservices practitioners’ suggested types of solutions in Table 5 and the Solution Taxonomy sheet in [replpack]. One representative quotation about the suggested solutions is depicted below.
@@
The Likert Distribution shows the level of agreement and importance for each issue, cause, and solution category. In contrast, the Mean indicates the average of the Likert distribution for the issue, cause, and solution categories. The p-value indicates statistical differences between Experience $\leq$ 6 Years and Experience > 6 Years in the Experience Based Grouping column. We used the non-parametric Mann–Whitney U test to test the null hypothesis (i.e., there is no significant difference between the responses in both groups). We describe the impact of the groups on the responses as significant if the p-value is less than 0.05 (see in Table 6). The Effect Size is measured by taking the mean difference of Experience > 6 Years and Experience $\leq$ 6 Years.
@@
4 Discussion and Needs for Future Research
In this section, we analyze the relationships between the identified issues, causes, and solutions (Section 4.1), and interpret these issue–cause–solution links derived from OSS issue discussions to explain dominance patterns and guide practitioners in prioritizing mitigation strategies. We then present the research implications (Section 4.2), which synthesize the study results, discuss implications for research and practice, and identify directions for future research (Section 4.3) on MSAs.
@@ 4.1 Relationship between issues, causes, solutions
While the taxonomy in Figure 4 provides a categorization of issues in microservices systems, a mapping between the issues, their causes, and solutions is presented in Figure 5. Mapping diagrams are frequently used in systematic mapping studies—relying on bubble plots—to correlate data or concepts along different dimensions [petersen2008systematic]. We have chosen the mapping diagram to depict the issues on the Y-axis and their causes on the X-axis, and to present the solutions that can address these issues at the intersections of the X and Y axes. Eight different types of symbols, labeled at the top of the figure, are used to represent various types of solutions. The interpretation of the mapping in Figure 5 involves locating a given issue on the Y-axis, associating this issue with its cause(s) on the X-axis, and identifying the possible solutions to address the issue, as elaborated and exemplified below. Figure 5 indicate that it is possible that one issue could cause other issues. Similarly, one cause may be the reason for several issues, or several causes may contribute to one issue.
@@
The mapping in Figure 5 can have a diverse interpretation based on the intent of the analysis thatămay include but is not limited to frequencyăanalysis and data correlations. It is virtually impossible to elaborate on all possible interpretations, however; to exemplify some of the possible interpretations can be as follows:
  • What are the most and least frequently occurring issues in microservices systems? As per the mapping, the most frequent (top 3) issues are related to Technical Debt, CI/CD, and Exception Handling, representing a total of 764 identified issues. On the other hand, the least frequent issues relate to Organizational, Update and Installation, and Typecasting issues. The dominance of Technical Debt and CI/CD in Figure 5 should be interpreted as a recurring pattern in OSS issue discussions, rather than as isolated implementation mistakes. In our issue–cause–solution links, Technical Debt frequently appears when changes propagate across multiple services and accumulate into Code Debt and Service Design Debt, while CI/CD reappears when pipelines, dependencies, and execution environments must continuously co-evolve across many services. This is consistent with the mapping evidence, where the dominant cause categories (e.g., General Programming Error, Invalid Configuration and Communication Problem, and Legacy Versions, Compatibility and Dependency Problem) repeatedly link to these issue categories, and where the associated fixes are mainly artifact-oriented (Fix Artifacts, Add Artifacts, and Modify Artifacts). In contrast, low-frequency issue categories (e.g., Organizational, Update and Installation, and Typecasting) should not be interpreted as negligible. Rather, they may be reported less consistently, absorbed into broader issue categories (e.g., upgrade problems reported under CI/CD/build or dependency issues), or discussed indirectly through symptoms that are recorded under higher-frequency issue categories. Therefore, frequency in our dataset captures what is most visible and repeatedly discussed in OSS issue trackers, while some low-frequency issue categories may still have high impact when they occur.
  • What are the most and least common causes of a specific category of issues? The mapping of the causes suggests that General Programming Errors, Invalid Configuration and Communication Problems along with Legacy Versions, Compatibility and Dependency Problem are the most common causes of the issues. Similarly, Fragile Code represents the least common cause for issues in microservices systems. A key interpretive insight from the cause distribution is that the most common causes are not narrowly tied to one development phase; instead, they represent recurring breakdowns in integration and evolution across services. General Programming Errors frequently occur because microservices multiply implementation points and boundary conditions, so local coding problems more often surface as system-level symptoms. Invalid Configuration and Communication Problems recur because microservices externalize behavior into configuration and networked interactions, where inconsistencies across services and environments can trigger runtime and CI/CD failures. Legacy Versions, Compatibility and Dependency Problem is common because microservices systems rely heavily on libraries, container images, and platform/tool versions that evolve asynchronously across services. Taken together, these dominant cause categories recur across multiple issue categories in Figure 5, which helps explain why Technical Debt, CI/CD, build, and service execution and communicating issues remain prominent in OSS issue discussions.
  • What are the most and least recurring solutions to fix a specific category of issues? Fixing Artifacts, Add Artifacts, and Modify Artifacts represent the most recurring solutions to address microservices issues, whereas importing/Export Artifacts, Upgrade Tools and Platforms, and Manage Configuration and Execution categories represent the least recurring solutions to address microservices issues. The distribution of solutions in Figure 5 suggests that OSS microservices maintenance is predominantly reactive and artifact-centric. Fixing Artifacts, Add Artifacts, and Modify Artifacts are the most recurring solutions because they are typically fast to apply, localized to specific code/configuration/build artifacts, and immediately verifiable through builds, tests, or pipeline reruns (e.g., patching source code, updating build scripts, adjusting configuration files, fixing deployment descriptors). In contrast, the least recurring solution categories (Import/Export Artifacts, Upgrade Tools and Platforms, and Manage Configuration and Execution) often require broader coordination across services and toolchains, can introduce new compatibility and dependency risks, and may affect multiple services simultaneously, which makes them harder to implement and validate in OSS settings. Overall, this gap indicates that while recurring issues are frequently resolved via local artifact updates, there is a need for research and tooling that better supports broader, system-level solutions (e.g., safer upgrade strategies and systematic configuration/execution management) to reduce repeated issue recurrence across microservices systems.
  • How practitioners can prioritize mitigation? Figure 5, together with the solution taxonomy (Table 4), supports an evidence-based prioritization that practitioners can apply as follows. First, prioritize the highest-frequency Issue Categories (RQ1) to focus effort where issues occur most often. Second, within those Issue Categories, target the most common Cause Categories (RQ2) because they recur across multiple Issue Categories and therefore represent recurring drivers of problems (e.g., General Programming Errors, Invalid Configuration and Communication Problems, and Legacy Versions, Compatibility and Dependency Problem). Third, choose an initial mitigation strategy using the most recurring Solution Categories (RQ3), such as Fixing Artifacts, Add Artifacts, and Modify Artifacts, and then complement these with less recurring but broader Solution Categories (e.g., Upgrade Tools and Platforms, Manage Configuration and Execution) when the mapping shows repeated cross-service occurrence. Overall, the issue–cause–solution links enable practitioners to decide what to address first (high-frequency Issue Categories), where to intervene (dominant Cause Categories), and how to proceed (recurring Solution Categories).